OpenStudy (anonymous):

which value of x is the solution of (x/ 3)+( x+1/2)= x

OpenStudy (anonymous):

its is x/3 not x3 my bad

OpenStudy (anonymous):

\[\frac{x}{3}+\frac{x+1}{2}=x\]

OpenStudy (anonymous):

ya there you go

OpenStudy (anonymous):

i need help with that